# Aggregation Models with Optimal Weights for Distributed Gaussian Processes

### v1

@@ 4.2 Distributed exact GP
The detailed procedures are outlined in Algorithm 5 and Algorithm 6 in Appendix. The time complexity of the parallelized Algorithm 6 is $\mathcal{O}(n_{i}^{3}+M^{3})$, where $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ for expert $\mathcal{M}_{i}$, and $M$ is the number of experts. A comparison of the time complexity of the aggregation models and the full exact GP is presented in the last row of Table 1.
5 Experiment
In this section, we evaluate the proposed algorithm against several baseline methods: full GP, PoE, gPoE, BCM, rBCM, and grBCM, using both synthetic data (Section 5.1) and UCI datasets (Asuncion et al.,, 2007) (Section 5.2). The size of the inducing variables is set to $m=128$. We use the radial basis function (RBF) kernel with separate lengthscales for each input dimension, defined by the following equation:
$k(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\exp\left(-\sum_{i=1}^{d}\frac{(\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i})^{2}}{2l_{i}^{2}}\right),$ (46)
where $\mathbf{x}_{i},\mathbf{x}^{\prime}_{i}\in\mathbb{R}$ are the $i$-th components of the inputs $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}$, $l_{i}$ is the lengthscale along the $i$-th dimension, and $\sigma_{f}^{2}>0$ is the signal variance.
@@ 5.1.1 Training
In this section, we assess the performance of four distributed training methods for SVGP: FACT, ADMM, FedAvg, and FedProx. We generate a training dataset of size $n=10^{4}$ and dimension $d=2$ using the RBF kernel with $(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f})=(0.3,1.3,0.1,0.2)$. The dataset is distributed among $M=2,4,8,19,20,40,80,100$ experts, respectively. We set the initial values of the hyperparameters $(l_{1}^{(0)},l_{2}^{(0)},\sigma_{\epsilon}^{(0)},\sigma_{f}^{(0)})=(2.0,0.3,1.0,\sqrt{1.5})$, and use Adam optimizer (Kingma and Ba,, 2014) with a learning rate of $0.1$ for all methods. For ADMM, the regularization parameter is set to $\rho=500$. For FedProx, the proximal coefficient is set to $\mu=0.01$, with a sampling rate of 0.5 for the local updates. The hyperparameters of the SVGP $\bm{\theta}=(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f},\mathbf{u})^{\top}$ are trained over 200 iterations, where $\mathbf{u}$ are the inducing variables of size $m=128$. The results are averaged over 10 different seeds for each setting of $M$.
Figure 2 illustrates the change of the parameters with the training iterations for different numbers of experts $M$. We observe that FACT and ADMM perform worse as the number of experts $M$ increases, while FedAvg performs the best overall, converging faster and more stably than FedProx. Figure 8 in Appendix shows boxplots of the hyperparameter estimates after 200 iterations, indicating that FedAvg converges closest to the ground truth with the least variability. Therefore, we use FedAvg for distributed training in the subsequent experiments with the UCI datasets.
@@ 5.1.2 Prediction
In this section, we evaluate the performance of aggregation models and full GP models without optimizing any hyperparameters. We use the Ackley function (Ackley,, 2012) defined over the interval $[-1,1]^{d}$, to benchmark the aggregation models:
$f(\mathbf{x})=-a\exp\left(-b\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\exp\left(\frac{1}{d}\sum_{i=1}^{d}\cos(cx_{i})\right)+a+\exp(1),\quad\mathbf{x}\in\mathbb{R}^{d},$ (47)
with parameters $a=20$, $b=0.2$, $c=2\pi$. We employ the RBF kernel with fixed hyperparameters $(l_{i},\sigma_{\epsilon}^{2},\sigma_{f}^{2})=(0.3,0.025,1.0)$ for all $i=1,\ldots,d$. For SVGP, we use $m=128$ inducing points, randomly generated with a fixed seed across all aggregation methods.
Figure 3: SVGP posteriors with the different aggregation models on the Ackley function on $n=10^{4}$ training points and $n_{t}=50$ test points with the number of experts $M=4$, dimension $d=1$, and the the number of inducing variables $m=128$. We denote the aggregated predictions (aggreg), by red lines, the full SVGP posteriors (full) by blue lines, and the ground truth values (true) by black dots.
Figure 3 shows the aggregated predictions of SVGP on $n=10^{4}$ training points and $n_{t}=50$ test points, with dimension $d=1$ and $M=4$ experts. The proposed method (opt) and five baselines (PoE, gPoE, BCM, rBCM, grBCM) are compared to the full SVGP. All methods use the same inducing points. It’s observed that PoE, gPoE, grBCM and the proposed method (opt) provide better predictive means than BCM and rBCM. However, the predictive variances of PoE, BCM, and rBCM shrink to zero and is not consistent with the full SVGP. Our method, the distributed SVGP with optimal weights, delivers accurate predictions for both the mean and the variance.
@@
Figure 4: Comparison of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. The number of experts considered are $M=2,4,8,10,20,40,80,100$. The lower $x$-axis represents the size of the local training dataset $n_{i}=n/M$, and the upper $x$-axis represents the number of the experts $M$. Left: Logarithm of time for computing the aggregated predictions. Middle: RMSE between the aggregated predictions and the ground truth. Right: NLPD between the aggregated predictions and the ground truth. Top: Mean of the computational time, RMSE and NLPD. Bottom: Standard deviation of the computational time, RMSE and NLPD.
@@
Figure 5: Comparison of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. The number of experts considered are $M=2,4,8,10,20,40,80,100$. The lower $x$-axis represents the size of the local training dataset $n_{i}=n/M$, and the upper $x$-axis represents the number of the experts $M$. Left: Logarithm of time for computing the aggregated predictions. Middle: RMSE between the aggregated predictions and the ground truth. Right: NLPD between the aggregated predictions and the ground truth. Top: Mean of the computational time, RMSE and NLPD. Bottom: Standard deviation of the computational time, RMSE and NLPD.
@@
Next, we evaluate the aggregation models on $n=10^{4}$ training data of dimension $d=2$ and $n_{t}=50\times 50=2500$ test points, evenly spaced over the interval $[-1,1]^{2}$. The training dataset is divided among $M=2,4,8,10,20,40,80,100$ experts, respectively. To eliminate the impact of lengthscales and seeds, we average all the metrics over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$ and 10 different seeds $1,2,3,4,5,6,7,8,9,10$ for each setting of $M$.

### v2

@@ 4.2 Distributed exact GP
The detailed procedures are outlined in Algorithm 5 and Algorithm 6. The time complexity of the parallelized Algorithm 6 is $\mathcal{O}(n_{i}^{3}+M^{3})$, where $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ for expert $\mathcal{M}_{i}$, and $M$ is the number of experts. A comparison of the time complexity of the aggregation models and the full exact GP is presented in the last row of Table 1.
4.3 Convergence
In this section, we introduce the theoretical properties of the proposed aggregation method with optimal weights. In contrast to PoE and BCM families whose weights are chosen heuristically, our weights are defined as the minimizer of a function space residual objective in eq. 21 and are computed from the resulting linear system in eq. 22. This optimization-based method is what enables the consistency for the aggregated predictor.
Setting.
Let the global dataset $\mathcal{D}$ be randomly partitioned into local datasets $\{\mathcal{D}_{i}\}_{i=1}^{M_{n}}$ without replacement, where the number of experts $M_{n}$ increases with the global dataset size $n$ such that (i) $\lim_{n\to\infty}M_{n}=\infty$, and (ii) $\liminf_{n\to\infty}\frac{n}{M_{n}^{2}}>0$. Then the local sample sizes satisfy $n_{i}\asymp n/M_{n}\to\infty$. Assume each expert $\mathcal{M}_{i}$ is a zero-mean GP with the same kernel and Gaussian noise model as the full GP.
For each GP expert, let $\mathcal{P}_{i}f$ denote the posterior operator conditioned on the local dataset $\mathcal{D}_{i}$ (as defined in eq. 21). Let $\mu_{i}(\cdot)$ and $\sigma_{i}^{2}(\cdot)$ be the predictive mean and variance of expert $i$, where $i=1,\ldots,M_{n}$, and let $\mu_{\text{full}}(\cdot)$ and $\sigma_{\text{full}}^{2}(\cdot)$ denote the predictive mean and variance of the full GP trained on $\mathcal{D}$.
Assumptions.
  1. 1.
    local predictive consistency.
    We assume a standard GP predictive consistency property: as the local sample size $n_{i}\to\infty$, the predictive mean and variance of the $i$-th expert converge to the same limiting predictive quantities as the full GP, at any fixed test input $\mathbf{x}_{*}$. This is a classical result in GP posterior consistency theory (Choi, 2005), and it is also the auxiliary fact from proof of Proposition 2 in (Liu et al., 2018) to control the asymptotic behavior of each expert’s predictive mean and variance. Specifically, for any fixed $\mathbf{x}_{*}$,
    $\lim_{n\to\infty}\max_{1\leq i\leq M_{n}}\big|\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big|=0,\qquad\lim_{n\to\infty}\max_{1\leq i\leq M_{n}}\big|\sigma_{i}^{2}(\mathbf{x}_{*})-\sigma_{\text{full}}^{2}(\mathbf{x}_{*})\big|=0.$ (27)
  2. 2.
    weight stability.
    We assume the optimizer $\bm{\beta}_{n}=(\beta_{1,n},\ldots,\beta_{M_{n},n})^{\top}$ returned by the linear system in eq. 22 is stable, i.e.:
    $\|\bm{\beta}_{n}\|_{\ell_{1}}\leq B\quad\text{for all sufficiently large $n$},$ (28)
    for some constant $B$ independent of $n$. This assumption is standard for linear estimators: without controlling the magnitude of $\bm{\beta}_{n}$, the residual term $\sum_{i=1}^{M_{n}}\beta_{i,n}(f-\mathcal{P}_{i}f)$ below may fail to vanish even if each $(f-\mathcal{P}_{i}f)\to 0$. In practice, such stability can be enforced by adding a small jitter regularizer to the Gram matrix in eq. 22.
Aggregation prediction.
The aggregated predictor of the proposed method has the form (see eqs. 24 and 26)
$\mu_{\mathcal{A}}(\mathbf{x}_{*})=\sum_{i=1}^{M_{n}}\beta_{i,n}\,\mu_{i}(\mathbf{x}_{*}),\qquad\sigma_{\mathcal{A}}^{2}(\mathbf{x}_{*})=\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\,\sigma_{i}^{2}(\mathbf{x}_{*}).$ (29)
Define the sum of weights $s_{n}:=\sum_{i=1}^{M_{n}}\beta_{i,n}$, then we have the decomposition as follows:
$\mu_{\mathcal{A}}(\mathbf{x}_{*})=s_{n}\,\mu_{\text{full}}(\mathbf{x}_{*})+\sum_{i=1}^{M_{n}}\beta_{i,n}\big(\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big).$ (30)
Thus, mean consistency reduces to proving (i) $s_{n}\to 1$, and (ii) the second term in eq. 30 converges to zero.
Mean.
  • Step 1: the objective value at the optimizer converges to zero.
    Define the function space residual objective from eq. 21
    $J_{n}(\bm{\beta})\;:=\;\Big\|f-\sum_{i=1}^{M_{n}}\beta_{i}\,\mathcal{P}_{i}f\Big\|.$ (31)
    Let $\bm{\beta}_{n}$ be a minimizer of $J_{n}(\bm{\beta})$. Then for any $j\in\{1,\dots,M_{n}\}$,
    $J_{n}(\bm{\beta}_{n})\;\leq\;J_{n}(\mathbf{e}_{j})\;=\;\|f-\mathcal{P}_{j}f\|,$ (32)
    where $\mathbf{e}_{j}$ is the $j$-th standard basis vector. By Assumption 1 and $n_{j}\asymp n/M_{n}\to\infty$, we have
    $\lim_{n\to\infty}\|f-\mathcal{P}_{j}f\|=0$ (33)
    Therefore,
    $\lim_{n\to\infty}J_{n}(\bm{\beta}_{n})=0$ (34)
  • Step 2: derive $s_{n}=\sum_{i=1}^{M_{n}}\beta_{i,n}\to 1$.
    Using the identity
    $f-\sum_{i=1}^{M_{n}}\beta_{i,n}\mathcal{P}_{i}f=(1-s_{n})f+\sum_{i=1}^{M_{n}}\beta_{i,n}\big(f-\mathcal{P}_{i}f\big),$ (35)
    we bound $|1-s_{n}|$ by applying the triangle inequality:
    $|1-s_{n}|\cdot\|f\|$ $=\big\|(1-s_{n})f\big\|$
    $\leq\Big\|f-\sum_{i=1}^{M_{n}}\beta_{i,n}\mathcal{P}_{i}f\Big\|+\Big\|\sum_{i=1}^{M_{n}}\beta_{i,n}(f-\mathcal{P}_{i}f)\Big\|$
    $\leq J_{n}(\bm{\beta}_{n})+\sum_{i=1}^{M_{n}}|\beta_{i,n}|\cdot\|f-\mathcal{P}_{i}f\|$
    $\leq J_{n}(\bm{\beta}_{n})+\|\bm{\beta}_{n}\|_{\ell_{1}}\,\max_{1\leq i\leq M_{n}}\|f-\mathcal{P}_{i}f\|.$ (36)
    By eq. 34, the first term $J_{n}(\bm{\beta}_{n})$ tends to 0. By Assumption 1, $\lim_{n\to\infty}\max_{i}\|f-\mathcal{P}_{i}f\|=0$, and by Assumption 2, $\|\bm{\beta}_{n}\|_{\ell_{1}}\leq B$. Hence the right hand side of section 4.3 converges to 0, which implies that
    $\lim_{n\to\infty}\sum_{i=1}^{M_{n}}\beta_{i,n}=\lim_{n\to\infty}s_{n}=1.$ (37)
  • Step 3: prove mean consistency. Return to eq. 30. The second term is bounded by
    $\Big|\sum_{i=1}^{M_{n}}\beta_{i,n}\big(\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big)\Big|\leq\|\bm{\beta}_{n}\|_{\ell_{1}}\,\max_{1\leq i\leq M_{n}}\big|\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big|\xrightarrow[]{}0\quad\text{as}\quad n\to\infty$ (38)
    using eqs. 28 and 27. Combining this with eq. 37 yields
    $\lim_{n\to\infty}\mu_{\mathcal{A}}(\mathbf{x}_{*})$ $=\lim_{n\to\infty}\Big(s_{n}\,\mu_{\text{full}}(\mathbf{x}_{*})+\sum_{i=1}^{M_{n}}\beta_{i,n}\big(\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big)\Big)$ (39)
    $=1\cdot\mu_{\text{full}}(\mathbf{x}_{*})+0$ (40)
    $=\mu_{\text{full}}(\mathbf{x}_{*}).$ (41)
    Therefore, the aggregated predictive mean is consistent.
Variance.
Using eq. 29 we get the following decomposition for aggregated predictive variance:
$\sigma_{\mathcal{A}}^{2}(\mathbf{x}_{*})$ $=\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\sigma_{i}^{2}(\mathbf{x}_{*})$
$=\Big(\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\Big)\sigma_{\text{full}}^{2}(\mathbf{x}_{*})+\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\big(\sigma_{i}^{2}(\mathbf{x}_{*})-\sigma_{\text{full}}^{2}(\mathbf{x}_{*})\big).$ (42)
If $\|\bm{\beta}_{n}\|_{\ell_{2}}$ is uniformly bounded, then the second term in section 4.3 goes to zero by Assumption 1. However, to obtain $\sigma_{\mathcal{A}}^{2}(\mathbf{x}_{*})\to\sigma_{\text{full}}^{2}(\mathbf{x}_{*})$ one further needs the stronger condition $\sum_{i}\beta_{i,n}^{2}\to 1$, which we do not enforce. Consequently, our theoretical guarantee is primarily for the predictive mean, while variance consistency would require additional constraints on $\bm{\beta}_{n}$.
Stability.
Assumption 2 is required for a rigorous argument, i.e., the term $\sum_{i}\beta_{i,n}(f-\mathcal{P}_{i}f)$ in eq. 35 may not vanish without controlling the magnitude of $\bm{\beta}_{n}$ even when each component $(f-\mathcal{P}_{i}f)\to 0$. In our method, $\bm{\beta}_{n}$ is obtained by solving the finite dimensional linear system eq. 22 derived from the first-order optimality conditions of the function space objective $J_{n}(\bm{\beta})$. The solution is stable whenever the corresponding Gram matrix is well-conditioned. In practice, we ensure numerical stability by adding a small jitter to eq. 22, which is standard in GP computations and typically yields bounded solutions.
5 Experiments
In this section, we evaluate the proposed algorithm against full GP and several baselines that aggregate predictions via weighted averaging: PoE, gPoE, BCM, rBCM, and grBCM, using both synthetic data (Section 5.1) and UCI datasets (Asuncion and Newman, 2007) (Section 5.2). We restrict baselines to aggregation methods because our contribution is an aggregation algorithm for distributed GP experts, scalable methods for centralized GP such as inducing points approximations, structured kernel interpolation, Vecchia framework, or stochastic dual descent solvers are complementary and can be used within each expert, but they do not provide a competing aggregation mechanism under the distributed learning setting. The size of the inducing variables is set to $m=128$. We use the radial basis function (RBF) kernel with separate lengthscales $l_{i}$ for each input dimension $i$, and $\sigma_{f}^{2}>0$ is the signal variance of the kernel. The code for the experiments is available at https://github.com/hchen19/optimal-distgp.
@@ 5.1.1 Training
In this section, we assess the performance of four distributed training methods for SVGP: FACT, ADMM, FedAvg, and FedProx. We generate a training dataset of size $n=10^{4}$ and dimension $d=2$ using the RBF kernel with $(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f})=(0.3,1.3,0.1,0.2)$. The dataset is distributed among $M=2,4,8,10,20,40,80,100$ experts, respectively. We set the initial values of the hyperparameters $(l_{1}^{(0)},l_{2}^{(0)},\sigma_{\epsilon}^{(0)},\sigma_{f}^{(0)})=(2.0,0.3,1.0,\sqrt{1.5})$, and use Adam optimizer (Kingma and Ba, 2015) with a learning rate of $0.1$ for all methods. For ADMM, the regularization parameter is set to $\rho=500$. For FedProx, the proximal coefficient is set to $\mu=0.01$, with a sampling rate of 0.5 for the local updates. The hyperparameters of the SVGP $\bm{\theta}=(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f},\mathbf{u})^{\top}$ are trained over 200 iterations, where $\mathbf{u}$ are the inducing variables of size $m=128$. The results are averaged over 10 different seeds for each setting of $M$.
Figure 2 illustrates the change of the parameters with the training iterations for different numbers of experts $M$. We observe that FACT and ADMM perform worse as the number of experts $M$ increases, while FedAvg performs the best overall, converging faster and more stably than FedProx. Figure 8 shows boxplots of the hyperparameter estimates after 200 iterations, indicating that FedAvg converges closest to the ground truth with the least variability. Therefore, we use FedAvg for distributed training in the experiments with the UCI datasets.
Figure 3: SVGP posteriors with the different aggregation models on $n=100$ training points in [1,1][-1,1] and $n_{t}=50$ test points in [1.2,1.2][-1.2,1.2] with the number of experts $M=4$, dimension $d=1$, and the number of inducing variables $m=128$. We denote the aggregated predictions (aggreg) by dashed lines, the full SVGP posteriors (full) by solid lines, and the observations by dots.
@@ 5.1.2 Prediction
In this section, we evaluate the performance of aggregation models and full GP models without optimizing any hyperparameters. We generate the ground truth dataset using RBF kernel with fixed hyperparameters $(l_{i},\sigma_{\epsilon}^{2},\sigma_{f}^{2})=(0.3,0.25,1.0)$ for all $i=1,\ldots,d$. For SVGP, we use $m=128$ inducing points, randomly generated with a fixed seed across all aggregation methods.
Figure 3 shows the aggregated predictions of SVGP on $n=100$ training points uniformly sampled in [1,1][-1,1] and $n_{t}=50$ test points evenly spaced in [1.2,1.2][-1.2,1.2], with dimension $d=1$ and $M=4$ experts. We compare our method (opt) against five baselines (PoE, gPoE, BCM, rBCM, grBCM) and the full SVGP model using the same inducing points. The proposed method yields predictions that closely match the full SVGP, especially in the extrapolation region outside the training domain (i.e., in $1.2,-1]\cup[1,1.2]$). In contrast, most baselines show degraded performance in these regions. Furthermore, the predictive variances of our method and grBCM align more closely with the full SVGP, suggesting better consistency of the predictions compared to other methods.
@@
Figure 4: Comparison of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points in $1,1]^{2}$ and $n_{t}=2500$ test points in $1.2,1.2]^{2}$. The number of experts considered are $M=2,4,8,10,20,40,80,100$. The lower $x$-axis represents the size of the local training dataset $n_{i}=n/M$, and the upper $x$-axis represents the number of the experts $M$. Left: Logarithm of time for computing the aggregated predictions. Middle Left: RMSE between the aggregated predictions and the ground truth. Middle Right: NLPD between the aggregated predictions and the ground truth. Right: 2-Wasserstein distance to the full GP. Top: Mean of the computational time, RMSE, NLPD, and 2-Wasserstein distance. Bottom: Standard deviation of the computational time, RMSE, NLPD, and 2-Wasserstein distance.
@@
Figure 5: Comparison of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points in $1,1]^{2}$, $m=128$ inducing points, and $n_{t}=2500$ test points in $1.2,1.2]^{2}$. The number of experts considered are $M=2,4,8,10,20,40,80,100$. The lower $x$-axis represents the size of the local training dataset $n_{i}=n/M$, and the upper $x$-axis represents the number of the experts $M$. Left: Logarithm of time for computing the aggregated predictions. Middle Left: RMSE between the aggregated predictions and the ground truth. Middle Right: NLPD between the aggregated predictions and the ground truth. Right: 2-Wasserstein distance to the full GP. Top: Mean of the computational time, RMSE, NLPD, and 2-Wasserstein distance. Bottom: Standard deviation of the computational time, RMSE, NLPD, and 2-Wasserstein distance.
@@
Next, we evaluate the aggregation models on $n=10^{4}$ training data of dimension $d=2$ in $1,1]^{2}$ and $n_{t}=50\times 50=2500$ test points, evenly spaced over the interval $1.2,1.2]^{2}$. The training dataset is divided among $M=2,4,8,10,20,40,80,100$ experts, respectively. To eliminate the impact of seeds, we average all the metrics over 10 different seeds for each setting of $M$.